Karsted Air Services is now in the final year of a project. The equipment originally cost $34 million, of which 80% has been depreciated. Karsted can sell the used equipment today for $8.5 million, and its tax rate is 40%. What is the equipment's after-tax salvage value? Round your answer to the nearest dollar. Write out your answer completely. For example, 13 million should be entered as 13,000,000. $ =?
Answer #1

Book value as on date of sale=Cost-Accumulated Depreciation

=34*(1-0.8)=6.8 million

Hence gain on sale=(8.5-6.8)=$1.7 million

Hence after-tax salvage value=Sale proceeds-(Tax rate*gain on sale)

=8.5-(0.4*1.7) million

=$7,820,000.